# Scribble-Supervised Medical Image Segmentation with Dynamic Teacher Switching and Hierarchical Consistency

### v1

@@ 1 Introduction
Despite their practicality, scribble annotations introduce severe supervision sparsity, providing only limited pixel-level information without clear boundary cues. This coarse supervision often leads to ambiguous pseudo-label propagation, causing the model to misinterpret unlabeled regions and generate inaccurate boundaries. As a result, networks tend to overfit annotated pixels and fail to generalize structural information to unlabeled areas. Prior works have sought to address this issue using pseudo-labeling or consistency regularization frameworks. For instance, Luo et al. [dmpls] proposed dynamically mixed pseudo-label supervision, Li et al. [scribblevc] introduced vision-class embeddings to enhance semantics, Chen et al. [ail] designed a cross-image matching strategy using a reference set constructed from class-specific tokens and pixel-level features, and Zhang et al. [helpnet] developed a hierarchical perturbation consistency strategy with an entropy-guided pseudo-label ensemble. Although these methods improved scribble-supervised segmentation, they still face two key issues: (1) Pseudo-label noise accumulation, where erroneous predictions are reinforced during self-training, and (2) Lack of adaptive supervision, as fixed teacher-student frameworks cannot consistently identify the most reliable guidance source.
@@ 2.1 Scribble-supervised learning
$\mathcal{L}_{Scribble}=\mathcal{L}_{pCE}(y_{S},y_{Scrib})$ (1)
@@ 2.2 Dynamic Teacher Switching
$\mathcal{L}_{T_{1}}=\mathcal{L}_{pCE}(y_{T_{1}},y_{Scrib}),\quad\mathcal{L}_{T_{2}}=\mathcal{L}_{pCE}(y_{T_{2}},y_{Scrib})$ (3)
@@
$y_{T}^{*}=\begin{cases}\text{Softmax}(y_{T_{1}}),&\text{if }\mathcal{L}_{T_{1}}<\mathcal{L}_{T_{2}}\\ \text{Softmax}(y_{T_{2}}),&\text{otherwise}\end{cases}$ (4)
@@
$y_{PL}=\begin{cases}\arg\max_{k}(y_{T}^{*})_{i},&\text{if }\max_{k}(y_{T}^{*})_{i}\geq\tau\\ \text{Ignore},&\text{otherwise}\end{cases}$ (5)
@@
$\mathcal{L}_{Pseudo}=\frac{1}{2}\Big(\mathcal{L}_{CE}(y_{S},y_{PL})+\mathcal{L}_{Dice}(y_{S},y_{PL})\Big)$ (6)
@@ 2.3 Hierarchical Consistency
$\mathcal{L}_{HiCo}=\frac{1}{2}\Big(\mathcal{L}_{feat}(F_{S}^{L},F_{T*}^{L})+\mathcal{L}_{feat}(F_{S}^{H},F_{T*}^{H})\Big)$ (8)
@@ 2.4 Total Loss Function
$\mathcal{L}_{Total}=\mathcal{L}_{Scribble}+\mathcal{L}_{Pseudo}+\mathcal{L}_{HiCo}$ (9)
@@
$\theta_{{T*}}^{t}=\alpha\theta_{{T*}}^{t-1}+(1-\alpha)\theta_{S}^{t}$ (10)

### v2

@@ 1 Introduction
Despite their practicality, scribble annotations introduce severe supervision sparsity, offering only limited pixel-level information without clear boundary cues. This coarse supervision often leads to ambiguous pseudo-label propagation, causing models to misinterpret unlabeled regions and produce inaccurate boundaries, ultimately resulting in overfitting to annotated pixels and poor generalization to unlabeled areas. Prior works have sought to mitigate these issues using pseudo-labeling or consistency regularization frameworks. For instance, Luo et al. [dmpls] proposed dynamically mixed pseudo-label supervision, Li et al. [scribblevc] introduced vision-class embeddings to enhance semantics, Chen et al. [ail] designed a cross-image matching strategy using a reference set constructed from class-specific tokens and pixel-level features, and Zhang et al. [helpnet] developed a hierarchical perturbation consistency strategy with an entropy-guided pseudo-label ensemble. Related advances in semi-supervised and source-free settings further highlight the importance of reliable pseudo-label selection and uncertainty-aware refinement [nguyen2026adaptive, nguyen2026up2d]. Although these methods improved scribble-supervised segmentation, they still face two key issues: (1) Pseudo-label noise accumulation, where erroneous predictions are reinforced during self-training, and (2) Lack of adaptive supervision, as fixed teacher-student frameworks cannot consistently identify the most reliable guidance source.
@@ 2.1 Scribble-supervised learning
$\mathcal{L}_{Scribble}=\mathcal{L}_{pCE}(y_{S},y_{Scrib}).$ (1)
@@ 2.2 Dynamic Teacher Switching
$\mathcal{L}_{T_{1}}=\mathcal{L}_{pCE}(y_{T_{1}},y_{Scrib}),\quad\mathcal{L}_{T_{2}}=\mathcal{L}_{pCE}(y_{T_{2}},y_{Scrib}).$ (3)
@@
$y_{T}^{*}=\begin{cases}\text{Softmax}(y_{T_{1}}),&\text{if }\mathcal{L}_{T_{1}}<\mathcal{L}_{T_{2}}\\ \text{Softmax}(y_{T_{2}}),&\text{otherwise}\end{cases}.$ (4)
@@
$y_{PL}=\begin{cases}\arg\max_{k}(y_{T}^{*})_{i},&\text{if }\max_{k}(y_{T}^{*})_{i}\geq\tau\\ \text{Ignore},&\text{otherwise,}\end{cases}$ (5)
@@
$\mathcal{L}_{Pseudo}=\frac{1}{2}\Big(\mathcal{L}_{CE}(y_{S},y_{PL})+\mathcal{L}_{Dice}(y_{S},y_{PL})\Big).$ (6)
@@ 2.3 Hierarchical Consistency
$\mathcal{L}_{HiCo}=\frac{1}{2}\Big(\mathcal{L}_{feat}(F_{S}^{L},F_{T*}^{L})+\mathcal{L}_{feat}(F_{S}^{H},F_{T*}^{H})\Big).$ (8)
@@ 2.4 Total Loss Function
$\mathcal{L}_{Total}=\mathcal{L}_{Scribble}+\mathcal{L}_{Pseudo}+\mathcal{L}_{HiCo},$ (9)
@@
$\theta_{{T*}}^{t}=\alpha\theta_{{T*}}^{t-1}+(1-\alpha)\theta_{S}^{t},$ (10)